# revtex4-2Repair the float

### v1

@@ Abstract
Moreover, there is a growing interest in preparing strongly correlated ultracold quantum bosonic gases within the quantum Hall regime [89, 90, 91], but achieving the necessary angular momentum poses a challenge. The rotation frequency required to generate such substantial angular momentum in a harmonically trapped BEC surpasses the deconfinement frequency. However, the ability to generate larger angular momentum in the bubble trap than in the harmonic trap makes it a promising candidate for obtaining a quantum Hall state. Furthermore, in experiments, the $\Delta$ parameter can be tuned by modulating the rf detuning. Increasing the value of $\Delta$ for a fixed rotation frequency $\Omega$ results in a larger moment of inertia of the condensate and facilitates access to larger angular momentum. Fig. 5(a), illustrates the growth of angular momentum with $\Delta$ for a constant rotation frequency $\Omega=0.9\omega$. In our numerical simulations, we vary the value of $\Delta$ from 00 (spherical harmonic trap) to $400l_{\rm{osc}}^{2}=243.4~{}\mu\rm{m^{2}}$ (bubble trap). For $\Delta=243.4~{}\mu\rm{m^{2}}$ and $\Omega=0.9\omega$, we observe that the angular momentum per particle reaches $276\hbar$. Further, due to the attractive DDI, the condensate’s chemical potential decreases, resulting in a decrease in sound velocity $v_{s}=\sqrt{\mu/m}$. Consequently, in this scenario, the local fluid velocity $v=\sqrt{\Delta_{\rm{max}}}\Omega\approx 8.817~{}\rm{mm/s}$ significantly surpasses the sound velocity $v_{s}\approx 0.124~{}\rm{mm/s}$, indicating that the condensate rotates at a hypersonic speed with a Mach number $v/v_{s}\approx 71$. Interestingly, as we increase $\Delta$, the transverse trapping frequency also rises (see Eq. (2) and Fig. 1), causing the rotation-induced droplets to merge with nearby ones and form macrodroplets immersed within a background superfluid [Fig. 5(a1)]. Further increasing the value of $\Delta$ results in the suppression of the density of the background superfluid, leading to the formation of an isolated macrodroplet crystal state [Fig. 5(a2)].
@@
Topological transition.- So far we have reported several intriguing phenomena that can emerge as a result of the unique bubble-shaped topology. In order to delve deeper into the role of the topology of the trap, we thoroughly investigate the impact of dynamical transitioning from a bubble-trapped BEC to a spherical-filled BEC and vice versa by tuning experimentally controlled parameter $\Delta$ with fixed $\Theta=0.0441l_{\rm{osc}}^{2}$.
@@
To initiate this study, we first consider a rotation-induced supersolid state in a bubble trap with $\Omega=0.53\omega$ and $\Delta=21.9~{}\mu\rm{m}^{2}$. Starting from this ground state configuration, we linearly decrease [^1] Irrespective of the initial state and trap rotation frequency, a sudden transition from a bubble trap to spherical harmonic trap results in the emergence of matter-wave interference and reduction in the condensate’s lifetime. See Fig. S4 in the Supplementary Material. the parameter $\Delta$ to a final value of $\Delta=0$, representing a spherical harmonic trap. This modulation of $\Delta$ is executed over a time period of $50~{}\rm{ms}$ and after which we let the condensate evolve for a certain time, keeping $\Omega$ constant. Initially, owing to the bubble-shaped trap geometry and influenced by rotation ($\Omega=0.53\omega$), we observe that the condensate exhibits a supersolid state [Fig. 6(a1)] and is spinning about the rotation axis of the bubble trap. This results in a relatively large angular momentum ($\langle\hat{L}_{z}\rangle/N\approx 18\hbar$) compared to the initial ground state [Fig. 6(b1)] in a harmonic trap ($\langle\hat{L}_{z}\rangle/N\approx 1.54\hbar$). During the linear decrease in the value of $\Delta$, the rotational symmetry ensures $[\hat{H},\hat{L}_{z}]=0$, implying that the expectation value of $\hat{L}_{z}$ remains conserved. Consequently, the angular momentum of the condensate, stemming from its rotation about the rotation axis in the bubble trap, is transferred to the condensate within the spherical harmonic trap in the form of topological excitations of vortices. Due to the significant angular momentum, the condensate’s radius expands and accommodates a large number of vortices distributed in a spiral pattern [see Figs. 6(a2) and 6(a3)].
As we stated, achieving a strongly correlated quantum Hall regime necessitates a large angular momentum, surpassing the deconfinement frequency limit for harmonically trapped ultracold gases. However, instead of following a direct dynamic stirring path, one can follow the abovementioned protocol. Starting from a rotating bubble trap with large $\Omega$ and $\Delta$, dynamically reduce the parameter $\Delta$ to zero. Given the ability to attain substantial angular momentum within the bubble trap and the conservation of angular momentum, this approach has the potential to facilitate the realization of the desired quantum Hall regime for a harmonically trapped BEC.
@@
We carry out the same dynamics in the opposite direction by changing the topology from a filled sphere ($\Delta=0$) to a hollow shell ($\Delta=21.9~{}\mu\rm{m}^{2}$). Initially, with the specified parameter value of $\epsilon_{\rm{dd}}=1.31$, the condensate within the harmonic trap exhibits a superfluid state. Under the influence of trap rotation ($\Omega=0.53\omega$), the condensate demonstrates four off-center vortices at $t=0~{}\rm{ms}$ [see Fig. 6(b1)], a characteristic feature of the superfluid phase. As we increase the value of $\Delta$, the density at the center diminishes and the vortices become nested at the center within a zero-density region. As a consequence of the conservation of angular momentum, the angular momentum of the vortices transforms into the spinning motion of the condensate about the rotation axis. Furthermore, analogous to the ground state of a bubble trap rotating with the same frequency ($\Omega=0.53\omega$) [see Fig. 3(d)], the increase in local DDI induces density modulation in the condensate, leading to the formation of a supersolid state [see Figs. 6(b2) and 6(b3)].
@@
Conclusions.- We have systematically examined the ground state phases of a dipolar BEC by varying the parameter $\epsilon_{\rm dd}$ within a bubble trap. Our investigation reveals that rotation applied to the bubble trap with certain ranges of rotation frequencies induces a transition from superfluid to supersolid phase, achieved through the modification of local DDI strength. Additionally, we have observed that condensate confined in a bubble trap can sustain higher circulation for an extended duration. Notably, our findings also demonstrate the emergence of macrodroplets rotating with hypersonic speed at higher rotation frequencies and larger values of $\Delta$. Furthermore, our analysis of the hollowing transition, a transition from a filled spherical BEC to a hollow BEC in the rotating frame, confirms that the condensate undergoes modulation due to the confinement imposed by the bubble trap. Another aspect of our study explores the topological change from a bubble trap to a filled BEC, demonstrating that tuning the trap parameter $\Delta$ can serve as a protocol to generate higher angular momentum states in harmonic confinement. This procedure has the potential to bring the interacting atomic gas into the desired quantum-Hall regime [92, 93]. The persistent current and hypersonic velocity exhibited by atomic dipolar BECs in a rotating bubble trap could pave the way for their utilization in quantum transport applications within atomtronics [94, 95, 96].
@@
Acknowledgment.- We thank Vito Scarola for carefully reading the manuscript and insightful discussions. We also thank Arpana Saboo for the helpful discussions. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of PARAM Shakti at IIT Kharagpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. H. S. Ghosh gratefully acknowledges the support from the Prime Minister’s Research Fellowship (PMRF), India. S. Halder acknowledges the MHRD Govt. of India for the research fellowship. S. Das acknowledges support from AFOSR FA9550-23-1-0034.

### v2

@@ IV Rotating ground state
Moreover, there is a growing interest in preparing strongly correlated ultracold quantum bosonic gases within the quantum Hall regime [102, 103, 104], but achieving the necessary angular momentum poses a challenge. The rotation frequency required to generate such substantial angular momentum in a harmonically trapped BEC surpasses the deconfinement frequency. However, the ability to generate larger angular momentum in the bubble trap than in the harmonic trap makes it a promising candidate for obtaining a quantum Hall state. Furthermore, in experiments, the $\Delta$ parameter can be tuned by modulating the rf detuning [68]. Increasing the value of $\Delta$ for a fixed rotation frequency $\Omega$ results in a larger moment of inertia of the condensate and facilitates access to higher angular momentum. Fig. 5(a), illustrates the growth of angular momentum with $\Delta$ for a constant rotation frequency $\Omega=0.9\omega$. In our numerical simulations, we vary the value of $\Delta$ from 00 (spherical harmonic trap) to $400l_{\rm{osc}}^{2}=243.4~{}\mu\rm{m^{2}}$ (bubble trap). For $\Delta=243.4~{}\mu\rm{m^{2}}$ and $\Omega=0.9\omega$, we observe that the angular momentum per particle reaches $276\hbar$ [105]. Further, due to the attractive DDI, the condensate’s chemical potential decreases, resulting in a decrease in sound velocity $v_{s}=\sqrt{\mu/m}$. Consequently, for $\Delta=21.9~{}\mu\rm{m^{2}}$ and $\Omega=0.9\omega$, the local fluid velocity $v=\sqrt{\Delta_{\rm{max}}}\Omega\approx 2.646~{}\rm{mm/s}$ significantly surpasses the sound velocity $v_{s}\approx 0.116~{}\rm{mm/s}$ 111 To assess the stability of the flow at this hypersonic speed, we perturb the system by an impurity. Notably, the flow remains stable against these perturbations, as illustrated in the Appendix B., indicating that the condensate rotates at a hypersonic speed with a Mach number $v/v_{s}\approx 22$ [78]. Interestingly, as we increase $\Delta$, the transverse trapping frequency also rises (see Eq. (2) and Fig. 1), causing the rotation-induced droplets to merge with nearby ones and form macrodroplets immersed within a background superfluid [Fig. 5(b)]. Further increasing the value of $\Delta$ results in the suppression of the density of the background superfluid, leading to the formation of an isolated macrodroplet crystal state [Fig. 5(c)] along with an increase of the Mach number. For $\Delta=243.4~{}\mu\rm{m^{2}}$ and $\Omega=0.9\omega$, the Mach number is close to $71$.
Figure 5: (a) Angular momentum $\langle\hat{L}_{z}\rangle/N\hbar$ of the ground state in the rotating frame shows linear variation with $\Delta$ for $\epsilon_{\rm dd}=1.31,\Omega=0.9\omega$ and $\Theta=0.0441l_{\rm{osc}}^{2}$. With these parameters, the inset (b) shows the density isosurfaces of a supersolid state at $\Delta=60.84~{}\mu\rm{m}^{2}$, and the inset (c) shows the density isosurfaces of a droplet state formed by macrodroplets at $\Delta=243.4~{}\mu\rm{m}^{2}$. Note that $\langle\hat{L}_{z}\rangle/N$ is approximate to $14\hbar$ for $\Delta=0$ and reaches close to $276\hbar$ at $\Delta=243.4~{}\mu\rm{m^{2}}$.
V Topological transition
@@
So far we have reported several intriguing phenomena that can emerge as a result of the unique bubble-shaped topology. In order to delve deeper into the role of the topology of the trap, we thoroughly investigate the impact of dynamical transitioning from a bubble-trapped BEC to a spherical-filled BEC and vice versa by tuning experimentally controlled parameter $\Delta$ with fixed $\Theta=0.0441l_{\rm{osc}}^{2}$.
@@
To initiate this study, we first consider a rotation-induced supersolid state in a bubble trap with $\Omega=0.53\omega$ and $\Delta=21.9~{}\mu\rm{m}^{2}$. Starting from this ground state configuration, we linearly decrease  [^2] Irrespective of the initial state and trap rotation frequency, a sudden transition from a bubble trap to a spherical harmonic trap results in the emergence of matter-wave interference and reduction in the condensate’s lifetime. See Fig. S4 in the Supplementary Material. the parameter $\Delta$ to a final value of $\Delta=0$, representing a spherical harmonic trap. This modulation of $\Delta$ is executed over a time period of $50~{}\rm{ms}$, after which we let the condensate evolve for a certain time, keeping $\Omega$ constant. Initially, owing to the bubble-shaped trap geometry and influenced by trap rotation at $\Omega=0.53\omega$, we observe that the condensate exhibits a supersolid state [Fig. 6(a1)] and is circulating about the rotation axis of the bubble trap. This results in a relatively high angular momentum ($\langle\hat{L}_{z}\rangle/N\approx 18\hbar$) compared to the ground state rotating with the same rotation frequency $\Omega=0.53\omega$ in a harmonic trap ($\langle\hat{L}_{z}\rangle/N\approx 1.54\hbar$) [Fig. 6(b1)]. During the linear decrease in the value of $\Delta$, the rotational symmetry ensures $[\hat{H},\hat{L}_{z}]=0$, implying that the expectation value of $\hat{L}_{z}$ remains conserved. Consequently, the angular momentum of the condensate, stemming from its rotation about the rotation axis in the bubble trap, is transferred to the condensate within the spherical harmonic trap in the form of topological excitations of vortices. Due to the significant angular momentum, the condensate’s radius expands and accommodates a large number of vortices distributed in a spiral pattern [see Figs. 6(a2) and 6(a3)].
@@
As we stated, achieving a strongly correlated quantum Hall regime necessitates a large angular momentum, surpassing the deconfinement frequency limit for harmonically trapped ultracold gases. However, instead of following a direct dynamic stirring path, one can follow the above-mentioned protocol. Starting from a rotating bubble trap with large $\Omega$ and $\Delta$, dynamically reduce the parameter $\Delta$ to zero. Given the ability to attain substantial angular momentum within the bubble trap and the conservation of angular momentum, this approach has the potential to facilitate the realization of the desired quantum Hall regime for a harmonically trapped BEC.
@@
We carry out the same dynamics in the opposite direction by changing the topology from a filled sphere ($\Delta=0$) to a hollow shell ($\Delta=21.9~{}\mu\rm{m}^{2}$). Initially, with the specified parameter value of $\epsilon_{\rm{dd}}=1.31$, the condensate within the harmonic trap exhibits a superfluid state. Under the influence of trap rotation ($\Omega=0.53\omega$), the condensate demonstrates four off-centered vortices at $t=0~{}\rm{ms}$ [see Fig. 6(b1)], a characteristic feature of the superfluid phase. As we increase the value of $\Delta$, the density at the center diminishes and the vortices become nested at the center within a zero-density region. As a consequence of the conservation of angular momentum, the angular momentum of the vortices transfers to the condensate, and the solid-body response of the density-modulated condensate causes its circulating motion about the rotation axis. Furthermore, analogous to the ground state of a bubble trap rotating with the same frequency ($\Omega=0.53\omega$) [see Fig. 3(d)], the increase in local DDI induces density modulation in the condensate, leading to the formation of a supersolid state [see Figs. 6(b2) and 6(b3)].
VI Conclusions and Outlook
@@
We have systematically examined the ground state phases of a dipolar BEC by varying the parameter $\epsilon_{\rm dd}$ within a bubble trap. Our investigation reveals that rotation applied to the bubble trap with certain ranges of rotation frequencies induces a transition from superfluid to supersolid phase, achieved through the modification of the local DDI strength. This transition to a supersolid state induced by rotation manifests as a first-order phase transition, while the transition resulting from tuning the scattering length exhibits a second-order nature. Additionally, we have observed that condensate confined in a bubble trap can sustain higher circulation for an extended duration. Notably, our findings also demonstrate the hypersonic flow, and the emergence of macrodroplets rotating with even higher Mach number at higher trap rotation frequencies and larger values of $\Delta$. Furthermore, our analysis of the hollowing transition in the rotating frame, confirms that the condensate undergoes modulation due to the confinement imposed by the bubble trap. Another aspect of our study explores the effect of topological change from a bubble trap to a filled BEC, demonstrating that tuning the trap parameter $\Delta$, can serve as a protocol to generate higher angular momentum states in harmonic confinement. This procedure has the potential to bring the interacting atomic gas into the desired quantum Hall regime [101, 100]. The persistent current and hypersonic velocity exhibited by atomic dipolar BECs in a rotating bubble trap could pave the way for their utilization in quantum transport applications within atomtronics [106, 107, 108].
Regarding the non-rotating case, we recently became aware of a very recent work [109] discussing the ground state phases of a dipolar BEC in a bubble trap.
Acknowledgments
@@
We thank Vito W. Scarola for carefully reading the manuscript and insightful discussions, and Uwe R. Fischer for their valuable comments and suggestions. We also thank Arpana Saboo for helpful discussions and carefully reading the manuscript. We acknowledge the National Supercomputing Mission for providing computing resources of PARAM Shakti at IIT Kharagpur. H. S. Ghosh gratefully acknowledges the support from the Prime Minister’s Research Fellowship (PMRF), India. S. Halder acknowledges the MHRD Govt. of India for the research fellowship. S. Das acknowledges support from AFOSR FA9550-23-1-0034.
APPENDIX A IMPACT OF ROTATION ON SUPERSOLID PHASE
In the main text, we have demonstrated the impact of rotation on the superfluid phase of a dipolar BEC confined in a bubble trap geometry. In this addendum, we explore the effect of rotation on the supersolid state. For this investigation, we begin with a non-rotating dipolar condensate confined in a bubble trap with $\Delta=21.9~{}\mu\rm{m}^{2}$, $\Theta=0.0441l_{\rm{osc}^{2}}$, and $\epsilon_{\rm{dd}}=1.37$. The ground state exhibits a supersolid state with regularly arranged ellipsoidal droplets along the trap and immersed in a background of low-density condensate with the superfluid fraction $f_{s}\approx 0.64$ as illustrated in Fig. 2(b). However, as we rotate the system, the density of background superfluid showcases sudden dips in its value at certain ranges of $\Omega$ values [see Figs. 7(a)-7(f)]. Consequently, the value of the superfluid fraction of the condensate drops at those
Figure 7: (a)-(f) Show the rotating ground state density profiles with $\epsilon_{\rm{dd}}=1.37$ for $\Omega=0.08\omega,0.20\omega,0.32\omega,0.52\omega,0.65\omega,$ and $0.78\omega$ in the $z=0$ plane, respectively. Other parameters are same as of Fig. 2(b).
specific ranges of $\Omega$ values [see Fig. 8]. This indicates that, under the influence of trap rotation, the supersolid state transitions towards an isolated droplet phase at certain ranges of rotation frequencies. Note that, in the droplet phase ($\epsilon_{\rm{dd}}>1.45$), the system only exhibits a solid-body response with the trap rotation, and the angular momentum changes linearly with $\Omega$.
APPENDIX B STABILITY OF THE FLOW
Figure 8: Superfluid fraction $f_{s}$ for rotating ground states at $\epsilon_{\rm{dd}}=1.37$ with trap rotation frequency $\Omega$. The dips at specific ranges of $\Omega$ values indicate the modification of the superfluid fraction. Other parameters are same as of Fig. 2(b).
We have observed that the condensate flows at a hypersonic speed in a rotating bubble trap with a large rotation frequency. For $\Delta=21.9~{}\mu\rm{m^{2}}$ and $\Omega=0.9\omega$, the condensate flows with a Mach number $v/v_{s}\approx 22$. Typically, according to the Landau criterion, above a critical velocity, excitation emerges in the system, leading to the breakdown of the frictionless flow. However, several studies have reported that when the kinetic energy of the system becomes sufficiently high to dominate all other energy scales, it suppresses excitations even above the Landau critical velocity [110, 111, 112, 113, 114, 115]. To illustrate the effect of perturbations on the stability of the flow, during time evolution we perturb the rotating ground state with $\epsilon_{\rm{dd}}=1.31$, $\Delta=21.9~{}\mu\rm{m}^{2}$ and $\Omega=0.9\omega$ by introducing an impurity potential of Gaussian form, defined as $V_{\rm{perturb}}=V_{0}\exp[-\{(x-x_{0})^{2}+(y-y_{0})^{2}+(z-z_{0})^{2}\}/\sigma^{2}]$, where $V_{0}$ is the potential height, $\sigma$ determines the width of the potential, and $\{x_{0},y_{0},z_{0}\}$ is the impurity position. We examine the stability of the flow against perturbations for various potential heights $V_{0}\approx\mu,~{}2\mu$ and $3\mu$, where $\mu$ is the chemical potential of the system. In each case, due to hypersonic velocity, the perturbations hardly affect the flow. Thus, the flow in our system is stabilized by its own angular momentum. Fig. 9 shows the time evolution of the perturbed system with $V_{0}\approx 3\mu$ and $\sigma=0.2l_{\rm{osc}}$.
Figure 9: Snapshots of the density distribution of the rotating ground state with $\epsilon_{\rm{dd}}=1.31$, $\Delta=21.9~{}\mu\rm{m}^{2}$ and $\Omega=0.9\omega$ in the $z=0$ plane, that is perturbed by a repulsive Gaussian potential with $V_{0}\approx 3\mu$, $\sigma=0.2l_{\rm{osc}}$, and $x_{0}=4.76~{}\mu\rm{m}$, $y_{0}=0$, $z_{0}=0$. The blue circular marker in (a1) and (a2) marks the position of the impurity.